# On De Giorgi’s Conjecture of Nonlocal approximations for free-discontinuity problems: The symmetric gradient case

### v1

@@ 3.1. Preliminaries
Therefore, for $\mathcal{L}^{n}$-a.e. $z\in\text{co}(\Sigma\cup\{0\})$ we find real smooth coefficients $\{c_{1}(z),\dotsc,c_{n}(z)\}$ such that $\phi^{+}(z)=\sum_{i}c_{i}(z)\phi^{+}(z-z_{i})$. In addition, combining (3.23) and (3.26) we infer for every $i=1,\dotsc,n$ that
$|(u_{r_{k}}(z_{i})-u_{r_{k}}(z))\cdot\phi^{+}(z-z_{i})|\to 0,\ \text{pointwisea.e. $z\in\text{co}(\Sigma\cup\{0\})$ as $k\to\infty$}.$
@@
$|u_{r_{k}}(z)\cdot\phi^{+}(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi^{+}(z-z_{i})|$ $\leq\sum_{i}|c_{i}(z)||(u_{r_{k}}(z)-u_{r_{k}}(z_{i}))\cdot\phi^{+}(z-z_{i})|$
$\leq\sum_{i}|c_{i}(z)||D(\hat{u}_{r_{k}})_{z_{i}}^{\phi^{+}(z-z_{i})}|(B_{1}(0)_{z_{i}}^{\phi^{+}(z-z_{i})})$
$\leq\sum_{i}|c_{i}(z)|g_{i,k}(\phi^{+}(z-z_{i})).$
By combining (3.24) and (3.25), the above inequality yields that for a.e. $z\in\text{co}(\Sigma\cup\{0\})$
$|u_{r_{k}}(z)\cdot\phi^{+}(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi^{+}(z-z_{i})|\to 0,\ \text{ as $k\to\infty$},$
thus, by recalling (3.22), we have for a.e. $z\in\text{co}(\Sigma\cup\{0\})$
$|f(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi^{+}(z-z_{i})|\to 0,\ \text{ as$k\to\infty$}.$
By performing the change of variables $z=t\xi$ for $(\xi,t)\in\Lambda^{+}\times\mathbb{R}\setminus\{0\}$, we rewrite the above convergence as: for $\mathcal{H}^{n-1}$-a.e. $\xi\in\Sigma$ it holds
$|a_{\xi}\,\text{sign}(t)+b_{\xi}-\sum_{i}c_{i}(t\xi)u_{r_{k}}(z_{i})\cdot\phi^{+}(t\xi-z_{i})|\to 0,\ \text{a.e. $t\in(-1,1)$ as $k\to\infty$}.$ (3.27)
We conclude by showing that (3.27) gives a contradiction. Indeed, fix a $\xi\in\Sigma$ for which (3.27) holds true. Consider the finite-dimensional real vector subspace $V$ of $L^{0}((-1,1))$ generated by the elements $\{v_{ij}\colon(-1,1)\to\mathbb{R}:i,j=1,\dotsc,n\}$, where $v_{ij}(t):=c_{i}(t\xi)\phi_{j}^{+}(t\xi-z_{i})$ ($\phi_{j}^{+}:=\phi^{+}\cdot e_{j}$). Condition (3.27) combined with Proposition 3.12 (recall that pointwise convergence implies convergence in measure) imply that the function $v\colon(-1,1)\to\mathbb{R}$, defined as $v(t):=a_{\xi}\,\text{sign}(t)+b_{\xi}$, belongs to $V$. Since the generators $v_{ij}$ are all smooth functions, any linear combination of them still belongs to $C^{\infty}((-1,1))$, forcing $a_{\xi}=0$. This is not possible because we assumed at the beginning $a_{\xi}\neq 0$ and we reach a contradiction. The proof is thus concluded. ∎
Before proving Theorem 3.6, we need to clarify the relation between the set appearing in formula (3.15) and the set where certain oscillations of $u$ do not vanish. In [2] the following notion of one-dimensional radial oscillation for a function $u\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ around $x\in\mathbb{R}^{n}$ was considered.
Definition 3.14 (One dimensional radial oscillation around $x$).
Let $f\colon\mathbb{R}\to\mathbb{R}$ be measurable. We introduce the oscillation of $f$ at scale $r>0$ around the origin as
$\text{Osc}_{r}(f):=\inf_{\text{Lip}(\theta)\leq 1}\int_{-1/4}^{1/4}(|f(rt)-\theta|\wedge 1)\,t^{n-1}\,\mathrm{d}t\,.$
For $\Omega\subseteq\mathbb{R}^{n}$ open and $u\colon\Omega\to\mathbb{R}^{n}$ measurable we define the one-dimensional radial oscillation of $u$ around $x\in\Omega$ as
$\text{Osc}(u,x):=\limsup_{r\searrow 0}\int_{\mathbb{S}^{n-1}}\text{Osc}_{r}(\hat{u}^{\xi}_{x})\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\,.$
The set of all points $x\in\Omega$ for which the one-dimensional radial oscillation of $u$ around $x$ do not vanish is given by
$\text{Osc}_{u}:=\{x\in\Omega:\text{Osc}(u,x)>0\}.$
We notice that, for every measurable function $u\colon\Omega\to\mathbb{R}^{n}$, the following inclusion holds true
$\{x\in\Omega:\,\text{$0\notin J_{\hat{u}^{\xi}_{x}}$ for $\mathcal{H}^{n-1}$-a.e.~{}$\xi\in\mathbb{S}^{n-1}$}\}\subset\text{Osc}_{u}.$
Therefore, proving that the set $\{x\in\Omega:\,0\notin J_{\hat{u}^{\xi}_{x}}\text{ for }\mathcal{H}^{n-1}\text{-a.e.~{}}\xi\in\mathbb{S}^{n-1}\}$ is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$ does not a priori imply the same for the set $\text{Osc}_{u}$, which corresponds to assumption (2) of [2, Theorem 4.12] with $\rho=1$. Nevertheless, in proving that the measure $\mathscr{I}_{u,1}$ is $(n-1)$-rectifiable, by virtue of Proposition 3.11 (which corresponds to [2, Proposition 4.7]), we can replace condition (2) of [2, Theorem 4.12] with
  1. (2’)
    The set $\{x\in\Omega:\,0\notin J_{\hat{u}^{\xi}_{x}}\text{ for }\mathcal{H}^{n-1}\text{-a.e.~{}}\xi\in\mathbb{S}^{n-1}\}$ is $\sigma$-finite w.r.t. $\mathcal{I}^{n-1}$,
where $\mathcal{I}^{n-1}$ denotes the $(n-1)$-dimensional integral geometric measure in $\mathbb{R}^{n}$.
@@
We start by showing that $\hat{\mu}_{u}^{p}(\Omega)<\infty$ implies $\mathscr{I}_{u,p}(\Omega)<\infty$. Indeed, given $\delta>0$, if we denote by $\mathscr{B}_{\delta,c}$ the set consisting of all countable covers of $\Omega$ made of open balls, we clearly have
$\mathscr{I}_{u,p}(\Omega)=\sup_{\delta>0}\inf_{G_{\delta}}\sum_{B\in G_{\delta}}\zeta_{p}(B)$ $\leq\sup_{\delta>0}\inf_{\mathscr{B}_{\delta,c}}\sum_{B\in\mathscr{B}_{\delta,c}}\bigg{(}\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}(B)^{p}\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\bigg{)}^{\frac{1}{p}}$
$\leq\sup_{\delta>0}\inf_{\mathscr{B}_{\delta,c}}\sum_{B\in\mathscr{B}_{\delta,c}}\hat{\mu}^{p}_{u}(B),$ (3.28)
where we used also that $\eta_{\xi}\leq\mu^{\xi}_{u}$ as measures for a.e. $\xi\in\mathbb{S}^{n-1}$, where $\eta_{\xi}$ is the measure in (3.13)–(3.14) associated to $u$, and that by construction
$\bigg{(}\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}(B)^{p}\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\bigg{)}^{\frac{1}{p}}\leq\hat{\mu}^{p}_{u}(B),\quad\text{for everyopen ball $B\subset\Omega$}.$
Finally, since $\hat{\mu}^{p}_{u}$ is assumed to be a finite (Borel) measure, then the outcome of (3.28) is exactly $\hat{\mu}^{p}_{u}(\Omega)$. This proves in particular the desired implication.
Now recall by [24] that given a Radon measure $\mu$, then the set $\{x:\Theta^{*n-1}(\mu,x)>0\}$ is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$. Therefore, we can combine Propositions 3.11 and 3.13 to infer that the assumption $\hat{\mu}_{u}^{p}(\Omega)<\infty$ implies that $\mathscr{I}_{u,1}$ is concentrated on a set which is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$. We are thus in position to apply [2, Theorem 4.12] with the condition (2) replaced by (2’) (see the discussion above) to infer exactly (3.5). The proof is thus concluded. ∎
@@
The proof follows the line of [21, Theorem 9.1], we report here only the main differences.
@@
As in [21, Theorem 9.1], the following parallelogram identity holds
@@
Since $e^{\xi}(x)$ is also positively homogeneous of degree $2$ by (3.31), we deduce by [20, Proposition 11.9] that for every $x\in\Omega\setminus N$ there exists a symmetric bilinear form $B_{x}:X\times X\to\mathbb{R}$ such that
@@ 4.1. Preliminary results
Let $p>1$, let $\Omega\subset\mathbb{R}^{n}$ be an open set, and let $\varphi_{p}\colon\mathbb{M}^{n\times n}\to[0,\infty)$ be the map with quadratic growth defined as
$\varphi_{p}(A)$ $:=\bigg{(}\int_{\mathbb{R}^{n}}\left(|A\xi\cdot\xi|^{2}\right)^{p}|\xi|^{p}e^{-|\xi|^{2}}\textnormal{d}\xi\bigg{)}^{\frac{1}{p}}\qquad\text{for$A\in\mathbb{M}^{n\times n}$ and $p<\infty$}.$
@@
Thanks to (4.19) and (4.32), it is not difficult to see that we are in position to apply Fréchet-Kolmogorov Theorem (see, e.g., [9, Theorem 4.26]) and obtain that the sequence $\{f_{k}\}_{k\in\mathbb{N}}$ is relatively compact in $L^{1}(F\times R)$ for every open set $F$ compactly contained in $\Omega$. Therefore, up to passing to a subsequence, we have $f_{k}\to f_{\infty}$ as $k\to\infty$ strongly in $L_{loc}^{1}(\Omega\times R)$. By a diagonal argument we can possibly pass to another subsequence such that
@@
Step 3: Let $\lambda>0$ and let $\tau_{\lambda}\colon\mathbb{R}\to[-\lambda,\lambda]$ be a strictly increasing function such that $\tau^{\prime}_{\lambda}\leq 1$, $\lim_{t\to\pm\infty}\tau_{\lambda}(t)=\pm\lambda$, and $\tau_{\lambda}(t)=t$ for $t\in[-\lambda/2,\lambda/2]$. Fix $\xi\in\mathbb{R}^{n}$, and let $y\in\Pi^{\xi}$. For every $\lambda>0$, we want to construct suitable modifications $\{v_{j,\lambda}\}_{j\in\mathbb{N}}\subset\text{SBV}(\Omega_{y}^{\xi})$ of $(\hat{u}_{k})_{y,\lambda}^{\xi}(t):=\tau_{\lambda}(u_{k}(y+t\xi)\cdot\xi)$, as in [26, Theorem 3.4, Step $2$], such that
@@
which implies that $u\in\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and $\hat{\mu}_{u}^{p}(\Omega)<\infty$. Thus by Theorem 3.1 we conclude that $u\in\text{GSBD}(\Omega)$. In addition, recall that, by Lemma 4.6, we have for every $c\in\mathbb{R}$
@@ 5.1. $\Gamma$-convergence
We start with the following pointwise convergence result following the strategy in [26, Theorem 3.4].
@@ 5.2. Convergence of quasi-minimisers
6. The case $p=1$
In this final section we recall and discuss the following problem which has already been formulated in the introduction.
Open Problem.
Let $u\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a measurable vector field, and define the Borel regular measure $\mathscr{I}_{u,1}$ in $\mathbb{R}^{n}$ as
@@ 6. The case $p=1$
If $u\in{\rm GBV}^{\mathcal{E}}(\mathbb{R}^{n})$, is it true that $\mathscr{I}_{u,1}$ is an $(n-1)$-rectifiable measure?
The consequence of an affirmative answer to the above problem is that for every $u\in{\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and for $\mathcal{H}^{n-1}$-a.e $\xi\in\mathbb{S}^{n-1}$
$(J_{u})^{\xi}_{y}=J_{\hat{u}^{\xi}_{y}},\quad\text{ for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$.}$ (6.1)
The key tool giving (6.1) from the above open problem, is a trace property for functions having bounded directional variation contained in [21, Theorem 5.1]. Loosely speaking, for a function $v\colon\Omega\to\mathbb{R}$ which is $L^{1}$-integrable and such that $D_{\xi}v$ is a measure, there exists a notion of traces $v^{\pm}$ a.e. on both sides of graphs which are given by Lipschitz functions of the form $g\colon\Pi^{\xi}\to\mathbb{R}$. Moreover, $v^{\pm}(y+t\xi)$ coincides with $v^{\xi\pm}_{y}(t)$ (up to a change of sign) for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$ and for every $t\in\text{graph}(g)^{\xi}_{y}$. This can be rigorously proved by following [21, Section 5].
Condition (6.1) is the key missing component needed to establish the desired $\Gamma$-convergence for $p=1$. In the language of Theorem 1.3, this is equivalent to proving that $\mathcal{I}^{s}_{u}=0$. This measure is finite and absolutely continuous with respect to the purely $(n-1)$-unrectifiable part of $\mathscr{I}_{u,1}$. Moreover, it assigns zero mass to any set that is $\sigma$-finite with respect to the $(n-1)$-dimensional Hausdorff measure. At the same time, it holds that $\mathcal{I}^{s}_{u}\,\bot\,\mathcal{L}^{n}$. While we believe the above open problem to be true, as suggested in [29], we cannot a priori exclude that $\mathscr{I}_{u,1}$ might concentrate on a purely $(n-1)$-unrectifiable set. Unfortunately, the techniques presented in this paper are insufficient to rule out this last scenario.
We dedicate this last part of the section to explain how Theorem 1.3 can be derived plus a few further comments. We observe that the results provided by Theorems 1.1 and 1.2 still hold for $p=1$ by replacing the space GSBD with $\textnormal{GSBV}^{\mathcal{E}}$ and the limit functional $\mathcal{F}^{p}$ with $\mathcal{F}$, this latter being defined as
$\mathcal{F}(u,\Omega):=\begin{cases}\int_{\mathbb{R}^{n}}\left(\int_{\Pi^{\xi}}\mathcal{MS}_{\frac{2}{\pi}}(\hat{u}_{y}^{\xi},\Omega_{y}^{\xi})\textnormal{d}\mathcal{H}^{n-1}(y)\right)|\xi|e^{-|\xi|^{2}}\textnormal{d}\xi&u\in\textnormal{GSBV}^{\mathcal{E}}{}(\Omega;\mathbb{R}^{n})\\ \infty&\text{otherwise in }L^{0}(\Omega;\mathbb{R}^{n}).\end{cases}$
By virtue of Theorem 3.7, a similar argument as in Lemma 4.6 gives for every $u\in\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$
$\int_{\mathbb{R}^{n}}\left(\int_{\Pi^{\xi}}\int_{\Omega_{y}^{\xi}}|\nabla\hat{u}_{y}^{\xi}|^{2}\,\textnormal{d}t\,\textnormal{d}\mathcal{H}^{n-1}(y)\right)|\xi|e^{-|\xi|^{2}}\textnormal{d}\xi=\frac{\pi^{\frac{n}{2}}}{2}\int_{\Omega}\Big{(}|e(u)|^{2}+\frac{1}{2}\textnormal{div}(u)^{2}\Big{)}\,\textnormal{d}x.$
Therefore, Theorem 1.3 follows by defining the measure $\mathcal{I}^{s}_{u}$ for every Borel set $B\subset\Omega$ as
$\mathcal{I}^{s}_{u}(B):=\int_{\mathbb{R}^{n}}\left(\int_{\Pi^{\xi}}\mathcal{H}^{0}(J_{\hat{u}_{y}^{\xi}}\cap B^{\xi}_{y})\textnormal{d}\mathcal{H}^{n-1}(y)\right)|\xi|e^{-|\xi|^{2}}\textnormal{d}\xi-\pi^{\frac{n-1}{2}}\mathcal{H}^{n-1}(B\cap J_{u}).$
In the computation of the $\Gamma$-liminf one can get rid of the measure $\mathcal{I}^{s}_{u}$. Indeed, by applying [21, Theorem 5.1] and since the jump set of a measurable function is always countably $(n-1)$-rectifiable (see [22]), the same aforementioned trace property for functions with directional bounded variation can be applied to show that every $u\in{\rm GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ satisfies for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ that
$(J_{u})^{\xi}_{y}\subset J_{\hat{u}^{\xi}_{y}},\quad\text{for $\mathcal{H}^{n-1}$-a.e. }y\in\Pi^{\xi}.$
Hence, we infer the following inequality for every $u\in\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$
$\int_{\mathbb{R}^{n}}\left(\int_{\Pi^{\xi}}\mathcal{H}^{0}(J_{\hat{u}_{y}^{\xi}})\textnormal{d}\mathcal{H}^{n-1}(y)\right)|\xi|e^{-|\xi|^{2}}\textnormal{d}\xi\geq\pi^{\frac{n-1}{2}}\mathcal{H}^{n-1}(J_{u})$ (6.2)
@@
$\frac{\pi^{\frac{n}{2}}}{2}\int_{\Omega}\Big{(}|e(u)|^{2}+\frac{1}{2}\textnormal{div}(u)^{2}\Big{)}\,\textnormal{d}x+\frac{\pi^{\frac{n+1}{2}}}{2}\mathcal{H}^{n-1}(J_{u}).$
The delicate part, which does not allow to conclude that $\mathcal{F}(u;\Omega)$ coincides with the above functional, is the construction of a recovery sequence. Here is exactly where the condition (6.1) becomes essential. Indeed, by following the proof of Theorem 1.2, the construction of a recovery sequence is immediately obtained by showing that (6.2) holds as an equality for every $u\in\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. As discussed earlier, this condition ultimately reduces to proving the rectifiability of $\mathscr{I}_{u,1}$.

### v2

@@ 3.1. Preliminaries
Therefore, for $\mathcal{H}^{n-1}$-a.e. $z\in\Pi^{\xi}\cap(B_{1}(0)\setminus B_{1/2}(0))$ we find real smooth coefficients $\{c_{1}(z),\dotsc,c_{n}(z)\}$ such that $\phi(z)=\sum_{i}c_{i}(z)\phi(z-z_{i})$. In addition, combining (3.25) and (3.27) we infer for every $i=1,\dotsc,n$ that
$|(u_{r_{k}}(z_{i})-u_{r_{k}}(z))\cdot\phi(z-z_{i})|\wedge 1\to 0,\ \text{pointwise for $\mathcal{H}^{n-1}$-a.e.~{}$z\in\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)}$ as $k\to\infty$}.$
@@
$|u_{r_{k}}(z)\cdot\phi(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi(z-z_{i})|\wedge 1$
$\leq\sum_{i}|c_{i}(z)||(u_{r_{k}}(z)-u_{r_{k}}(z_{i}))\cdot\phi(z-z_{i})|\wedge 1$
$\leq\sum_{i}|c_{i}(z)||\hat{\mu}_{u_{r}}^{\phi(z-z_{i})}|(B_{1}(0))$
By (3.26), the above inequality yields that for $\mathcal{H}^{n-1}$-a.e. $z\in\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)}$
$|u_{r_{k}}(z)\cdot\phi(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi(z-z_{i})|\to 0,\ \text{ as $k\to\infty$}.$
Thus, by recalling (3.24), we have for $\mathcal{H}^{n-1}$-a.e. $z\in\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)}$
$|f_{\nu,\xi}(z)-\sum_{i}c_{i}(z)u_{r_{k}}(z_{i})\cdot\phi(z-z_{i})|\to 0,\ \text{ as $k\to\infty$}.$ (3.28)
We conclude by showing that (3.28) gives a contradiction. Consider the finite-dimensional real vector subspace $V$ of $L^{0}\big{(}\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)};\mathbb{R}^{n}\big{)}$ generated by the elements $\{v_{ij}\colon\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)}\to\mathbb{R}^{n}:i,j=1,\dotsc,n\}$, where $v_{ij}(z):=c_{i}(z)\phi_{j}(z-z_{i})$ ($\phi_{j}:=\phi\cdot e_{j}$). Condition (3.28) combined with Proposition 3.15 (recall that pointwise convergence implies convergence in measure) implies that the function $f_{\nu,\xi}$ restricted to $\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)}$ belongs to $V$. Since the generators $v_{ij}$ are all smooth functions, any linear combination of them still belongs to $C^{\infty}(\Pi^{\xi}\cap\big{(}B_{1}(0)\setminus B_{\frac{1}{2}}(0)\big{)};\mathbb{R}^{n})$, forcing $a_{\nu,\xi}=0$. This is not possible because we assumed at the beginning $a_{\nu,\xi}\neq 0$ and we reach a contradiction. The proof is thus concluded. ∎
@@
We proceed by induction. Observe that in dimension one the theorem is trivially true. Now, assume that the theorem holds true in dimension $n-1$. Therefore, from Proposition 3.14 we infer that the measure $\mathscr{I}_{u,1}$ concentrates on the complement in $\Omega$ of the set $E$ defined in (3.18). Since by Proposition 3.16 we have that $\Omega\setminus E\subset\{x\in\Omega:\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)>0\}$, then by applying the classical density estimates for Radon measure (cf. [29]) we infer that $\Omega\setminus E$ is $\sigma$-finite with respect to $\mathcal{H}^{n-1}$. We are thus in position to apply Besicovitch-Federer structure theorem [29] to write $\Omega\setminus E=R\cup U$ where $R\subset\Omega$ is countably $(n-1)$-rectifiable while the Borel set $U\subset\Omega$ satisfies
$\mathcal{H}^{n-1}(\pi_{\xi}(U))=0,\ \ \text{ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$}.$
Eventually, from the very construction of $\mathscr{I}_{u,1}$, the above property immediately implies that $\mathscr{I}_{u,1}(U)=0$ and hence $\mathscr{I}_{u,1}$ concentrates on the rectifiable set $R$. In particular we deduce the fundamental inclusion
$J_{\hat{u}^{\xi}_{y}}\subset R^{\xi}_{y},\ \ \text{ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$}.$
We can then conclude by arguing as in the proof of [2, Theorem 4.12].
@@
The proof follows the line of [26, Theorem 9.1], we report here only the main differences.
@@
As in [26, Theorem 9.1], the following parallelogram identity holds
@@
Since $e^{\xi}(x)$ is also positively homogeneous of degree $2$ by (3.31), we deduce by [25, Proposition 11.9] that for every $x\in\Omega\setminus N$ there exists a symmetric bilinear form $B_{x}:X\times X\to\mathbb{R}$ such that
@@ 4.1. Preliminary results
Let $p\geq 1$, let $\Omega\subset\mathbb{R}^{n}$ be an open set, and let $\varphi_{p}\colon\mathbb{M}^{n\times n}_{sym}\to[0,\infty)$ be the map with quadratic growth defined as
$\varphi_{p}(A)$ $:=\bigg{(}\int_{\mathbb{R}^{n}}\left(|A\xi\cdot\xi|^{2}\right)^{p}|\xi|^{p}e^{-|\xi|^{2}}\textnormal{d}\xi\bigg{)}^{\frac{1}{p}}\qquad\text{for$A\in\mathbb{M}^{n\times n}_{sym}$ and $p<\infty$}.$
@@
Thanks to (4.19) and (4.32), it is not difficult to see that we are in position to apply Fréchet-Kolmogorov Theorem (see, e.g., [11, Theorem 4.26]) and obtain that the sequence $\{f_{k}\}_{k\in\mathbb{N}}$ is relatively compact in $L^{1}(F\times R)$ for every open set $F$ compactly contained in $\Omega$. Therefore, up to passing to a subsequence, we have $f_{k}\to f_{\infty}$ as $k\to\infty$ strongly in $L_{loc}^{1}(\Omega\times R)$. By a diagonal argument we can possibly pass to another subsequence such that
@@
Step 3: Let $\lambda>0$ and let $\tau_{\lambda}\colon\mathbb{R}\to[-\lambda,\lambda]$ be a strictly increasing function such that $\tau^{\prime}_{\lambda}\leq 1$, $\lim_{t\to\pm\infty}\tau_{\lambda}(t)=\pm\lambda$, and $\tau_{\lambda}(t)=t$ for $t\in[-\lambda/2,\lambda/2]$. Fix $\xi\in\mathbb{R}^{n}$, and let $y\in\Pi^{\xi}$. For every $\lambda>0$, we want to construct suitable modifications $\{v_{j,\lambda}\}_{j\in\mathbb{N}}\subset\text{SBV}(\Omega_{y}^{\xi})$ of $(\hat{u}_{k})_{y,\lambda}^{\xi}(t):=\tau_{\lambda}(u_{k}(y+t\xi)\cdot\xi)$, as in [31, Theorem 3.4, Step $2$], such that
@@
which implies that $\hat{\mu}_{u}^{p}(\Omega)<\infty$, and hence, by Remark 2.9, also that $u\in\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. Thus by Theorem 3.1 we conclude that $u\in\text{GSBD}(\Omega)$. In addition, recall that, by Lemma 4.6, we have for every $c\in\mathbb{R}$
@@ 5.1. $\Gamma$-convergence
We start with the following pointwise convergence result following the strategy in [31, Theorem 3.4].
@@ 5.2. Convergence of quasi-minimisers
Appendix A
Proof of Proposition 3.15.
@@
By renormalization, with no loss of generality, we may assume that for every $i=1,\dotsc,k$ and every $j=1,2,\dotsc$ it holds true $|\alpha^{j}_{i}|\leq 1$ and there exists at least one $i(j)=1,\dotsc,k$ such that $|\alpha^{j}_{i(j)}|=1$. Up to pass to a not relabelled subsequence in $j$, we may suppose that the index $i(j)$ does not depend on $j$, and that $\alpha^{j}_{i}\to\alpha_{i}$ for every $i=1,\dotsc,n$, for some real coefficients $\{\alpha_{1},\dotsc,\alpha_{k}\}$ which are not all identically zero. Therefore, since $K_{j}\nearrow A$ in measure we deduce that $\sum_{i=1}^{k}\alpha^{j}_{i}i_{K_{j}}(v_{i})$ (extended to zero out of $K_{j}$) converges to $\sum_{i=1}^{k}\alpha_{i}v_{i}$ pointwise a.e. in $A$ as $j\to\infty$. From (A.1) we immediately deduce that $\sum_{i=1}^{k}\alpha_{i}v_{i}=0$ in $L^{0}(A)$ which gives a contradiction. The claim is thus proved.